After spending 2/5 of her money on sweets and 1/6 on cosmetics, a woman was left with $6.50. How much money did she have originally?
Answer by rfer(16322)

30/30-1/6-2/5=
30/30-5/30-12/30=13/30
6.50/(13/30)=
6.50*30/13=$15